OpenStudy (anonymous):

Help Please! Solve the inequality. Show your work. 6y – 8 ≤ 10 +8 +8 6y 18 6y/6 = 0 18/6 = 3 y=3 The figures above are similar. Find the missing length. Show your work. plug the 3 into y place 6(3) =18 -8 = 10 18- 8 ≤ 10 10 ≤ 10

OpenStudy (anonymous):

6y/6 = y not 0

OpenStudy (anonymous):

Oops thanks

OpenStudy (anonymous):

What do you need help with? Everything is correct

OpenStudy (anonymous):

The figures above are similar. Find the missing length. Show your work. plug the 3 into y place 6(3) =18 -8 = 10 18- 8 ≤ 10 10 ≤ 10

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

The y was plugged in correctly

OpenStudy (anonymous):

Which figures?

OpenStudy (anonymous):

The figures above are similar. Find the missing length. Show your work. is the missing length 10?

OpenStudy (anonymous):

No.. I don't understand what the inequality has to do w/ the missing figure. Can you provide pics?

OpenStudy (anonymous):

If it were based on that inequality, the missing length would be LESS THAN OR EQUAL TO 10.

OpenStudy (anonymous):

It doesn't have pics

OpenStudy (anonymous):

So i just need to reverse my greater than or equal to to less than or equal too.

OpenStudy (anonymous):

No.. it's the same

OpenStudy (anonymous):

So what did i mess up on?

OpenStudy (anonymous):

You didn't mess up on anything. That is what I've been saying the entire time. Your Q entirely doesn't make sense because the Q states "The figures above are similar"

OpenStudy (anonymous):

So there needs to be some figures

OpenStudy (anonymous):

Okay, there isn't any picture with the work i was giving thanks though.